one volume of a compound of carbon , hydrogen and oxygen was exploded with 2.5 volumes of oxygen .the resultant volume contained 2 volumes of water vapour and 2 volumes of c02. all volumes were measured in identical conditions. determine the formula of the compound
 

Grade:11

1 Answers

abhay bajaj
6 years ago
Let compound be CxHyOz Therfore balanced eqn is CxHyOz +(x +y /4-z /2) - - - - - >xCO2 +y/2 H2O Therfore V Litres CxHyOz will react with(x +y/4-z/2) L O2 to form x)L CO2 and y/2L H2OFrm the question ;V(x +y /4-z /2)=2.5Vx=2And Vy/2=2Solving these :Vz=1;Vx=2;Vy=4Mole fraction of H:V(y) /V(x+y+z) =4/7Mole fraction of C:V(x) /V(x+y+z) =2/7Mole fraction of O:V(z) /V(x+y+z) =1 /7Hence compound is C2H4O OR CH3CHOHope it helped youThank you!!!!